# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ Abstract
Observations of gravitational-wave signals emitted by compact binary inspirals provide unique insights into their properties, but their analysis requires accurate and efficient waveform models. Intermediate- and extreme-mass-ratio inspirals (I/EMRIs), with mass ratios $q\gtrsim 10^{2}$, are promising sources for future detectors such as the Laser Interferometer Space Antenna (LISA). Modelling waveforms for these asymmetric-mass binaries is challenging, entailing the tracking of many harmonic modes over thousands to millions of cycles. The FastEMRIWaveforms (few) modelling framework addresses this need, leveraging precomputation of mode data and interpolation to rapidly compute adiabatic waveforms for eccentric inspirals into zero-spin black holes. In this work, we extend few to model eccentric equatorial inspirals into black holes with spin magnitudes $|a|\leq 0.999$. Our model supports eccentricities $e\leq 0.9$ and semi-latus recta $p\leq 200$, enabling the generation of long-duration IMRI waveforms, and produces waveforms in $\sim 100\,\mathrm{ms}$ with hardware acceleration. Characterising systematic errors, we estimate that our model attains mismatches of $\sim 10^{-5}$ (for LISA sensitivity) with respect to error-free adiabatic waveforms over the majority of the parameter space. We find that kludge models can introduce errors in signal-to-noise ratios (SNRs) as great as ${}^{+60\%}_{-40\%}$ and induce marginal biases of up to $\sim 1\,\sigma$ in parameter estimation. We show that LISA’s horizon redshift for I/EMRI signals varies significantly with $a$, reaching a redshift of $3$ ($15$) for EMRIs (IMRIs) with only minor $(\sim 10\%)$ dependence on $e$ for an SNR threshold of 20. For signals with $\rm{SNR}\sim 50$, spin and eccentricity-at-plunge are measured with uncertainties of $\delta a\sim 10^{-7}$ and $\delta e_{\rm f}\sim 10^{-5}$. This work advances the state-of-the-art in waveform generation for asymmetric-mass binaries, providing open-source tools for the investigation of I/EMRI astrophysics and data analysis.
@@ I Introduction
The advancement of gravitational wave (GW) astronomy has significantly enhanced our understanding of astrophysical phenomena, particularly through the analysis of compact binary mergers identified by ground-based LIGO-Virgo-KAGRA GW detector network KAGRA:2021vkt ; LIGOScientific:2014pky ; VIRGO:2014yos ; KAGRA:2018plz ; KAGRA:2021duu . To facilitate these efforts, models for the gravitational waveforms emitted by these coalescences have been developed and refined over time Varma:2019csw ; Ramos-Buades:2023ehm ; Pratten:2020ceb . However, there exists a region in the parameter space that has not yet been explored by GW detectors: systems where the masses $m_{1,2}$ of the binary components differ substantially (i.e., $m_{1}\gg m_{2}$onsortiumWaveformWorkingGroup:2023arg . These asymmetric-mass binaries are not the primary focus of current ground-based detectors, which have identified sources with mass ratios of up to $q=m_{1}/m_{2}\sim 10$ (in the case of GW190814 LIGOScientific:2020zkf ; LIGOScientific:2024elc ; LIGOScientific:2020stg ), and even up to $q\sim 26$ in the case of the event GW191219_163120 KAGRA:2021vkt (although this mass ratio estimate may be unreliable due to the inaccuracy of the waveform models used in this region of parameter space).
@@
Achieving these scientific goals entails the unbiased analysis of these signals, a procedure reliant upon accurate waveform models. Waveforms from IMRIs and EMRIs are among the most challenging to construct accurately of any compact binary system pound2022black ; barack2018self ; barack2009gravitational . As the trajectory of the secondary object is (in general) expected to be eccentric and inclined with respect to the primary object’s spin-momentum vector Alexander:2017aln (leading to Lense–Thirring precession of the orbital plane), the inspiral is tri-periodic Schmidt:2002qk , which manifests as a rich structure of tens of thousands of strong harmonic (sideband) modes in the gravitational waveform Hughes:2021exa ; Drasco:2005kz (as shown in Fig. 1). Models must accurately track the GW phase evolution of these modes to sub-radian precision over the thousands to millions of orbital cycles that the secondary object typically completes in-band, which is an infeasible task for the numerical relativity (NR) techniques that underpin existing models of comparable-mass binaries (despite ongoing efforts to address scale disparity in NR simulations Lousto:2020tnb ; Rosato:2021jsq ; Lousto:2022hoq ; Wittek:2024gxn ; Wittek:2024pis ). Moreover, waveform models must also be highly efficient; computational wall-times of less than one second are vital for the analysis of IMRIs and EMRIs to be a feasible prospect, especially in the context of a global LISA analysis framework (e.g., Refs. Katz:2024oqg ; Deng:2025wgk ; Littenberg:2023xpl ).
@@
The most promising approach that satisfies these stringent modelling constraints is the gravitational self-force (GSF) paradigm Mino:1996nk ; Quinn:1996am ; gralla2008rigorous ; Detweiler:2011tt ; Gralla:2012db ; pound2012second ; Harte:2014wya , where the metric of the binary is found by perturbatively solving the Einstein Field Equations in powers of the small mass ratio around the metric of the primary black hole; we refer readers to Refs. Poisson:2011nh ; barack2018self ; pound2022black for reviews. In particular, the modern waveform generation scheme in the GSF paradigm is built upon the multiscale framework Hinderer:2008dm ; Miller:2020bft ; pound2022black ; Mathews:2025nyb , leveraging the quasi-periodic orbital dynamics of asymmetric-mass binaries (see, e.g, Refs. Yunes:2010zj ; Chiaramello:2020ehz ; Zhang:2020rxy ; Ramos-Buades:2021adz ; Zhang:2021fgy ; Albanesi:2021rby ; Nagar:2022fep ; Albertini:2022dmc ; Albertini:2023aol ; Albanesi:2023bgi ; Albertini:2023aol ; Cheung:2023lnj ; Gamboa:2024hli ; Loutrel:2024qxp ; Paul:2024ujx ; Cheung:2024byb ; Akpinar:2025huz ; Morras:2025nlp ; Planas:2025feq ; Strusberg:2025qfv for other efforts towards IMRIs and EMRIs waveform generation). An important feature of this framework is that it enables waveform generation to be divided into an expensive “offline” step (in which many numerical data products are computed once and stored) and a fast “online” step in which these data products are interpolated and integrated to build waveforms quickly and accurately. In the case of the inspiral dynamics [^1] In this work, we exclusively focus on the inspiral stage of waveform modelling, as the merger-ringdown stage does not significantly contribute to the measurement precision of EMRI parameters. While our waveform model does not incorporate these stages of coalescence, they may become more impactful for IMRIs and their inclusion is worth investigating in this regard. Recent advancements have been made in terms of merger-ringdown modelling in the GSF framework (e.g. Refs. apte2019exciting ; Kuchler:2024esj ; Becker:2024xdi ; Lhost:2024jmw ; Kuchler:2025hwx ), highlighting the potential for the future extension of our model to include this stage; see also our concluding discussion in Section VI., at leading order, one can approximate the change of energy, angular momentum, and Carter constant Carter:1968rr of the binary due to the emission of GWs using so-called “flux-balance formulae” Mino:2003yg ; Drasco:2005is ; Sago:2005fn ; Tanaka:2005ue ; Isoyama:2018sib ; Grant:2024ivt in what is known as the adiabatic approximation Hughes:2001jr ; Finn:2000sy ; Mino:2006em ; Sundararajan:2008zm ; Hughes:2021exa ; Isoyama:2021jjd ; Kerachian:2023oiw ; Nasipak:2023kuf . These fluxes have been calculated using both numerical codes Nakamura:1987zz ; Shibata:1993uk ; Shibata:1993yf ; Shibata:1994xk ; kennefick1998stability ; Hughes:1999bq ; Glampedakis:2002ya ; Drasco:2005kz ; Hughes:2005qb ; Mino:2005an ; Fujita:2009us ; Harms:2014dqa ; Fujita:2020zxe ; Chen:2023lsa , and in analytic post-Newtonian (PN) expansions Mino:1997bx ; Sasaki:2003xr ; Ganz:2007rf ; Fujita:2012cm ; Shah:2014tka ; Fujita:2014eta ; Munna:2020iju ; Munna:2020som ; Munna:2023vds ; Sago:2024mgh ; Castillo:2024isq . It has been determined through a multiscale analysis (and subsequently verified with data analysis simulations in Ref. Burke:2023lno ) that one must go beyond leading order adiabatic waveforms and include the first post-adiabatic corrections to the dynamics of the binary to achieve the sub-radian accuracy required for LISA’s scientific goals. This requires knowledge of not only the first order in $\epsilon=1/q$ GSF Barack:2007tm ; barack2009gravitational ; Barack:2010tm ; Osburn:2014hoa ; Isoyama:2014mja ; Wardell:2015kea ; vandeMeent:2016pee ; vandeMeent:2017bcc ; Dolan:2023enf but also the second order GSF Pound:2019lzj ; Warburton:2021kwk ; Durkan:2022fvm  [^2] The effects of the secondary object’s spin also enter at this first post-adiabatic order Drummond:2023wqc ; Skoupy:2023lih ; Piovano:2024yks ; Mathews:2025nyb .. When included, GSF waveforms show remarkable agreement with NR even at $q\sim 10$ Wardell:2021fyy . These corrections can be included in an efficient way by leveraging averaging techniques Tanaka:2005ue ; Mino:2005an ; Mino:2006em , the multi-scale expansion Hinderer:2008dm ; Miller:2020bft ; pound2022black ; Mathews:2025nyb or near-identity averaging transformations VanDeMeent:2018cgn ; Lynch:2021ogr ; Lynch2022 ; Lynch:2023gpu ; Drummond:2023wqc ; Lynch:2024ohd . This leads to the expensive offline step mentioned above, where one densely tiles the intrinsic EMRI parameter space such that it may be rapidly interpolated during waveform generation. As the GSF data varies smoothly over parameter space, techniques such as cubic spline interpolation can be applied at low computational cost, facilitating rapid modelling of the inspiral dynamics and trajectory evolution. Unlike the dynamics, the amplitude evolution of each waveform mode needs only to be known to leading order for EMRIs (see, e.g., Appendix C of Ref. Burke:2023lno ). These amplitudes also vary smoothly over parameter space and are readily interpolated.
@@
Figure 1: Waveform for a retrograde eccentric EMRI into a spinning MBH in the frequency (upper-left), time-frequency (upper-right) and time domain (early and late times in lower-left and lower-right respectively). This system has parameters $\{m_{1},m_{2},a,p_{0},e_{0},d_{\mathrm{L}}\}=\{10^{5}\,M_{\odot},30\,M_{\odot},-0.998,28.3,0.85,1\,\mathrm{Gpc}\}$, and plunges after one year. When observed with LISA over this duration, this signal has an SNR of 41. The rich harmonic mode structure of the waveform evolves as the binary circularizes and the trajectory enters the strong-field region of the MBH (see also Fig. 4). Despite the waveform’s complexity and size, it is generated by few in less than $100\,\mathrm{ms}$ of wall-time for an A100 GPU, sufficiently fast enough for full parameter estimation studies to be performed on a timescale of hours.
@@
In this work, we address this need with the release of few v2, extending its domain of validity to spinning systems and covering a large portion of the parameter space of Kerr eccentric equatorial inspirals at adiabatic order. Specifically, we support dimensionless primary spin magnitudes of up to $0.999$ that exceeds the Thorne limit ($\sim 0.998$ Thorne:1974ve ) which represents the astrophysical “limit" of Kerr-spin parameters accounting for spin-up (and spin-down) mechanisms arising from accretion [^3] Methods for spinning up MBHs beyond the Thorne limit are discussed in bardeen1970kerr ; S_dowski_2011 ; Arbey_2020 ; further investigations regarding EMRIs into these near-extremal black holes can be found in gralla2016inspiral ; Compere:2017hsi ; burke2020transition ; burke2020constraining .. Similar to expanding coverage of parameter space in primary spin, we also extend our coverage of semi-latus rectum and eccentricity to $\sim 200$ and $0.9$, respectively (c.f., $\sim 16$ and $0.7$ in few v1). This provides the necessary support for the generation of long-duration IMRI waveforms. We also take this opportunity to thoroughly assess the systematics of waveform generation with few (with specific focus on the impact of interpolation errors), ensuring that our model produces waveforms that meet the accuracy requirements imposed by LISA (when neglecting post-adiabatic effects). From the results of this assessment, we conclude that our model is most robust for spins of up to $\sim 0.998$ and initial eccentricities of up to $\sim 0.85$. While our model does not include the post-adiabatic corrections necessary for the unbiased analysis of real signals (see Ref. Burke:2023lno for a discussion on this point), these additional contributions may be readily folded in as they become available, and it reproduces many features of these signals (such as detectability and parameter estimation precision) sufficiently well for use in studies of IMRI and EMRI science. As our model also accurately produces the rich harmonic mode spectrum of these sources (Fig. 1), it will also serve as an effective test-bed for the development of search and identification algorithms for these signals (which is an area of active research Cole:2025sqo ; Strub:2025dfs ; Ye:2023lok ; chua2022one ; Badger:2024rld ; Zhang:2022xuq ; Yun:2023vwa ) and will facilitate the construction of realistic test datasets as part of upcoming LISA Data Challenges Baghi:2022ucj .
@@
This paper is organized as follows. In Section II, we provide an overview of the few framework for fast waveform generation. We then describe the extensions to this framework (both in the incorporation of primary spin and in the general improvement of the software) that define our waveform model in Section III. To verify that our model is accurate and robust, we perform an extensive validation study of our systematic assumptions in Section IV. After assessing the computational cost of our model, we apply it in the examination of multiple facets of EMRI science in Section V. Our conclusions and outlook of future work are then presented in Section VI.
@@
Some incidental matters are relegated to Appendices. We explain the new mass convention that few v2 adopts in Appendix A. The data grid and coordinates that we use for the model implementation are explained in Appendix B. The source parameter prior probability distribution for the Monte-Carlo studies in this work is summarized in Appendix C, and we provide a concise description of our GW data analysis methods (which we apply throughout this work) in Appendix D. In Appendix E, we provide additional validation of our waveform model by comparing it to existing models and data from a specific cases. Finally, the marginal posterior distributions corresponding to the analyses performed in Section V.6 are provided in Appendix F.

### v2

@@ Abstract
Observations of gravitational-wave signals emitted by compact binary inspirals provide unique insights into their properties, but their analysis requires accurate and efficient waveform models. Intermediate- and extreme-mass-ratio inspirals (I/EMRIs), with mass ratios $q\gtrsim 10^{2}$, are promising sources for future detectors such as the Laser Interferometer Space Antenna (LISA). Modelling waveforms for these asymmetric-mass binaries is challenging, entailing the tracking of many harmonic modes over thousands to millions of cycles. The FastEMRIWaveforms (few) modelling framework addresses this need, leveraging precomputation of mode data and interpolation to rapidly compute adiabatic waveforms for eccentric inspirals into zero-spin black holes. In this work, we extend few to model eccentric equatorial inspirals into black holes with spin magnitudes $|a|\leq 0.999$. Our model supports eccentricities $e\leq 0.9$ and semi-latus recta $p\leq 200$, enabling the generation of long-duration IMRI waveforms, and produces waveforms in $\sim 100\,\mathrm{ms}$ with hardware acceleration. Characterising systematic errors, we estimate that our model attains mismatches of $\sim 10^{-5}$ (for LISA sensitivity) with respect to error-free adiabatic waveforms over the majority of the parameter space. We find that kludge models can introduce errors in signal-to-noise ratios (SNRs) as great as $^{+60\%}_{-40\%}$ and induce marginal biases of up to $\sim 1\,\sigma$ in parameter estimation. We show that LISA’s horizon redshift for I/EMRI signals varies significantly with $a$, reaching a redshift of 33 ($15$) for EMRIs (IMRIs) with only minor $(\sim 10\%)$ dependence on $e$ for an SNR threshold of 20. For signals with $\rm{SNR}\sim 50$, spin and eccentricity-at-plunge are measured with uncertainties of $\delta a\sim 10^{-7}$ and $\delta e_{\rm f}\sim 10^{-5}$. This work advances the state-of-the-art in waveform generation for asymmetric-mass binaries, providing open-source tools for the investigation of I/EMRI astrophysics and data analysis.
@@ I Introduction
The advancement of gravitational wave (GW) astronomy has significantly enhanced our understanding of astrophysical phenomena, particularly through the analysis of compact binary mergers identified by ground-based LIGO-Virgo-KAGRA GW detector network KAGRA:2021vkt ; LIGOScientific:2014pky ; VIRGO:2014yos ; KAGRA:2018plz ; KAGRA:2021duu . To facilitate these efforts, models for the gravitational waveforms emitted by these coalescences have been developed and refined over time Varma:2019csw ; Ramos-Buades:2023ehm ; Pratten:2020ceb . However, there exists a region in the parameter space that has not yet been explored by GW detectors: systems where the masses $m_{1,2}$ of the binary components differ substantially (i.e., $m_{1}\gg m_{2}$) LISAConsortiumWaveformWorkingGroup:2023arg . These asymmetric-mass binaries are not the primary focus of current ground-based detectors, which have identified sources with mass ratios of up to $q=m_{1}/m_{2}\sim 10$ (in the case of GW190814 LIGOScientific:2020zkf ; LIGOScientific:2024elc ; LIGOScientific:2020stg ), and even up to $q\sim 26$ in the case of the event GW191219_163120 KAGRA:2021vkt (although this mass ratio estimate may be unreliable due to the inaccuracy of the waveform models used in this region of parameter space).
@@
Achieving these scientific goals entails the unbiased analysis of these signals, a procedure reliant upon accurate waveform models. Waveforms from IMRIs and EMRIs are among the most challenging to construct accurately of any compact binary system pound2022black ; barack2018self ; barack2009gravitational . As the trajectory of the secondary object is (in general) expected to be eccentric and inclined with respect to the primary object’s spin-momentum vector Alexander:2017aln (leading to Lense–Thirring precession of the orbital plane), the inspiral is tri-periodic Schmidt:2002qk , which manifests as a rich structure of tens of thousands of strong harmonic (sideband) modes in the gravitational waveform Hughes:2021exa ; Drasco:2005kz (as shown in Fig.˜1). Models must accurately track the GW phase evolution of these modes to sub-radian precision over the thousands to millions of orbital cycles that the secondary object typically completes in-band, which is an infeasible task for the numerical relativity (NR) techniques that underpin existing models of comparable-mass binaries (despite ongoing efforts to address scale disparity in NR simulations Lousto:2020tnb ; Rosato:2021jsq ; Lousto:2022hoq ; Wittek:2024gxn ; Wittek:2024pis ). Moreover, waveform models must also be highly efficient; computational wall-times of less than one second are vital for the analysis of IMRIs and EMRIs to be a feasible prospect, especially in the context of a global LISA analysis framework (e.g., Refs. Katz:2024oqg ; Deng:2025wgk ; Littenberg:2023xpl ).
@@
The most promising approach that satisfies these stringent modelling constraints is the gravitational self-force (GSF) paradigm Mino:1996nk ; Quinn:1996am ; gralla2008rigorous ; Detweiler:2011tt ; Gralla:2012db ; pound2012second ; Harte:2014wya , where the metric of the binary is found by perturbatively solving the Einstein Field Equations in powers of the small mass ratio around the metric of the primary black hole; we refer readers to Refs. Poisson:2011nh ; barack2018self ; pound2022black for reviews. In particular, the modern waveform generation scheme in the GSF paradigm is built upon the multiscale framework Hinderer:2008dm ; Miller:2020bft ; pound2022black ; Mathews:2025nyb , leveraging the quasi-periodic orbital dynamics of asymmetric-mass binaries (see, e.g, Refs. Yunes:2010zj ; Chiaramello:2020ehz ; Zhang:2020rxy ; Ramos-Buades:2021adz ; Zhang:2021fgy ; Albanesi:2021rby ; Nagar:2022fep ; Albertini:2022dmc ; Albertini:2023aol ; Albanesi:2023bgi ; Albertini:2023aol ; Cheung:2023lnj ; Gamboa:2024hli ; Loutrel:2024qxp ; Paul:2024ujx ; Cheung:2024byb ; Akpinar:2025huz ; Morras:2025nlp ; Planas:2025feq ; Strusberg:2025qfv for other efforts towards IMRIs and EMRIs waveform generation). An important feature of this framework is that it enables waveform generation to be divided into an expensive “offline” step (in which many numerical data products are computed once and stored) and a fast “online” step in which these data products are interpolated and integrated to build waveforms quickly and accurately. In the case of the inspiral dynamics [^1] In this work, we exclusively focus on the inspiral stage of waveform modelling, as the merger-ringdown stage does not significantly contribute to the measurement precision of EMRI parameters. While our waveform model does not incorporate these stages of coalescence, they may become more impactful for IMRIs and their inclusion is worth investigating in this regard. Recent advancements have been made in terms of merger-ringdown modelling in the GSF framework (e.g. Refs. Apte:2019txp ; Kuchler:2024esj ; Becker:2024xdi ; Lhost:2024jmw ; Kuchler:2025hwx ), highlighting the potential for the future extension of our model to include this stage; see also our concluding discussion in Section VI., at leading order, one can approximate the change of energy, angular momentum, and Carter constant Carter:1968rr of the binary due to the emission of GWs using so-called “flux-balance formulae” Mino:2003yg ; Drasco:2005is ; Sago:2005fn ; Tanaka:2005ue ; Isoyama:2018sib ; Grant:2024ivt in what is known as the adiabatic approximation Hughes:2001jr ; Finn:2000sy ; Mino:2006em ; Sundararajan:2008zm ; Hughes:2021exa ; Isoyama:2021jjd ; Kerachian:2023oiw ; Nasipak:2023kuf . These fluxes have been calculated using both numerical codes Nakamura:1987zz ; Shibata:1993uk ; Shibata:1993yf ; Shibata:1994xk ; kennefick1998stability ; Hughes:1999bq ; Glampedakis:2002ya ; Drasco:2005kz ; Hughes:2005qb ; Mino:2005an ; Fujita:2009us ; Harms:2014dqa ; Fujita:2020zxe ; Chen:2023lsa , and in analytic post-Newtonian (PN) expansions Mino:1997bx ; Sasaki:2003xr ; Ganz:2007rf ; Fujita:2012cm ; Shah:2014tka ; Fujita:2014eta ; Munna:2020iju ; Munna:2020som ; Munna:2023vds ; Sago:2024mgh ; Castillo:2024isq . It has been determined through a multiscale analysis (and subsequently verified with data analysis simulations in Ref. Burke:2023lno ) that one must go beyond leading order adiabatic waveforms and include the first post-adiabatic corrections to the dynamics of the binary to achieve the sub-radian accuracy required for LISA’s scientific goals. This requires knowledge of not only the first order in $\epsilon=1/q$ GSF Barack:2007tm ; barack2009gravitational ; Barack:2010tm ; Osburn:2014hoa ; Isoyama:2014mja ; Wardell:2015kea ; vandeMeent:2016pee ; vandeMeent:2017bcc ; Dolan:2023enf but also the second order GSF Pound:2019lzj ; Warburton:2021kwk ; Durkan:2022fvm  [^2] The effects of the secondary object’s spin also enter at this first post-adiabatic order Drummond:2023wqc ; Skoupy:2023lih ; Piovano:2024yks ; Mathews:2025nyb .. When included, GSF waveforms show remarkable agreement with NR even at $q\sim 10$ Wardell:2021fyy . These corrections can be included in an efficient way by leveraging averaging techniques Tanaka:2005ue ; Mino:2005an ; Mino:2006em , the multi-scale expansion Hinderer:2008dm ; Miller:2020bft ; pound2022black ; Mathews:2025nyb or near-identity averaging transformations VanDeMeent:2018cgn ; Lynch:2021ogr ; Lynch2022 ; Lynch:2023gpu ; Drummond:2023wqc ; Lynch:2024ohd . This leads to the expensive offline step mentioned above, where one densely tiles the intrinsic EMRI parameter space such that it may be rapidly interpolated during waveform generation. As the GSF data varies smoothly over parameter space, techniques such as cubic spline interpolation can be applied at low computational cost, facilitating rapid modelling of the inspiral dynamics and trajectory evolution. Unlike the dynamics, the amplitude evolution of each waveform mode needs only to be known to leading order for EMRIs (see, e.g., Appendix C of Ref. Burke:2023lno ). These amplitudes also vary smoothly over parameter space and are readily interpolated.
@@
Figure 1: Waveform for a retrograde eccentric EMRI into a spinning MBH in the frequency (upper-left), time-frequency (upper-right) and time domain (early and late times in lower-left and lower-right respectively). This system has parameters $\{m_{1},m_{2},a,p_{0},e_{0},d_{\mathrm{L}}\}=\{10^{5}\,M_{\odot},30\,M_{\odot},-0.998,28.3,0.85,1\,\mathrm{Gpc}\}$, and plunges after one year. When observed with LISA over this duration, this signal has an SNR of 41. The rich harmonic mode structure of the waveform evolves as the binary circularizes and the trajectory enters the strong-field region of the MBH (see also Fig.˜4). Despite the waveform’s complexity and size, it is generated by few in less than $100\,\mathrm{ms}$ of wall-time for an A100 GPU, sufficiently fast enough for full parameter estimation studies to be performed on a timescale of hours.
@@
In this work, we address this need with the release of few v2, extending its domain of validity to spinning systems and covering a large portion of the parameter space of Kerr eccentric equatorial inspirals at adiabatic order. Specifically, we support dimensionless primary spin magnitudes of up to $0.999$ that exceeds the Thorne limit ($\sim 0.998$ Thorne:1974ve ) which represents the astrophysical “limit" of Kerr-spin parameters accounting for spin-up (and spin-down) mechanisms arising from accretion [^3] Methods for spinning up MBHs beyond the Thorne limit are discussed in bardeen1970kerr ; S_dowski_2011 ; Arbey_2020 ; further investigations regarding EMRIs into these near-extremal black holes can be found in gralla2016inspiral ; Compere:2017hsi ; burke2020transition ; burke2020constraining .. Similar to expanding coverage of parameter space in primary spin, we also extend our coverage of semi-latus rectum and eccentricity to $\sim 200$ and $0.9$, respectively (c.f., $\sim 16$ and $0.7$ in few v1). This provides the necessary support for the generation of long-duration IMRI waveforms. We also take this opportunity to thoroughly assess the systematics of waveform generation with few (with specific focus on the impact of interpolation errors), ensuring that our model produces waveforms that meet the accuracy requirements imposed by LISA (when neglecting post-adiabatic effects). From the results of this assessment, we conclude that our model is most robust for spins of up to $\sim 0.998$ and initial eccentricities of up to $\sim 0.85$. While our model does not include the post-adiabatic corrections necessary for the unbiased analysis of real signals (see Ref. Burke:2023lno for a discussion on this point), these additional contributions may be readily folded in as they become available, and it reproduces many features of these signals (such as detectability and parameter estimation precision) sufficiently well for use in studies of IMRI and EMRI science. As our model also accurately produces the rich harmonic mode spectrum of these sources (Fig.˜1), it will also serve as an effective test-bed for the development of search and identification algorithms for these signals (which is an area of active research Cole:2025sqo ; Strub:2025dfs ; Ye:2023lok ; chua2022one ; Badger:2024rld ; Zhang:2022xuq ; Yun:2023vwa ) and will facilitate the construction of realistic test datasets as part of upcoming LISA Data Challenges Baghi:2022ucj .
@@
This paper is organized as follows. In Section˜II, we provide an overview of the few framework for fast waveform generation. We then describe the extensions to this framework (both in the incorporation of primary spin and in the general improvement of the software) that define our waveform model in Section˜III. To verify that our model is accurate and robust, we perform an extensive validation study of our systematic assumptions in Section˜IV. After assessing the computational cost of our model, we apply it in the examination of multiple facets of EMRI science in Section˜V. Our conclusions and outlook of future work are then presented in Section˜VI.
@@
Some incidental matters are relegated to Appendices. We explain the new mass convention that few v2 adopts in Appendix˜A. The data grid and coordinates that we use for the model implementation are explained in Appendix˜B. The source parameter prior probability distribution for the Monte-Carlo studies in this work is summarized in Appendix˜C, and we provide a concise description of our GW data analysis methods (which we apply throughout this work) in Appendix˜D. In Appendix˜E, we provide additional validation of our waveform model by comparing it to existing models and data from a specific cases. Finally, the marginal posterior distributions corresponding to the analyses performed in Section˜V.6 are provided in Appendix˜F.